# Federated Instruction Tuning of LLMs with Domain Coverage Augmentation

### v5

@@ C.5 Augmentation Strategy Visualization
To more intuitively compare the domain coverage of different instruction augmentation methods, we randomly sample 5,000 instructions obtained through these methods and 10,000 in-domain instructions as the background, representing the distribution of specific domains in the public dataset. We then visualized the results using t-SNE van der Maaten and Hinton [2008], as shown in Figure 10. The plot shows that FedDCA encompasses most of the in-domain data, which is consistent with FedDCA’s domain coverage of each domain shown in Table 2. Also, we can observe that the random sampling strategy selects a lot of out-of-domain data while does not have good coverage in specific domains.
(a) Code
(b) Medical
(c) Financial
(d) Mathematical
@@
Figure 10: Visualization of cross-client data distribution in different domains, performing t-SNE dimensionality reduction on retrieved instructions through various augmentation strategies. We randomly sample 10,000 in-domain samples as background while randomly sampling 5,000 samples from the cross-client augmented dataset for different instruction augmentation methods for comparison.

### v6

@@ Appendix D Augmentation Strategy Visualization
To more intuitively compare the domain coverage of different instruction augmentation methods, we randomly sample 5k cross-client instructions obtained through these methods and 10k in-domain instructions as the background, representing the distribution of specific domains in the public dataset. We then visualized the results using t-SNE van der Maaten and Hinton (2008), as shown in Figure 10. The plot shows that FedDCA encompasses most of the in-domain data, which is consistent with FedDCA’s domain coverage of each domain shown in Table 1. Also, we can observe that the random sampling strategy selects a lot of out-of-domain data while does not have good coverage in specific domains.
(a) Medical
(b) Financial
(c) Mathematical
@@
Figure 10: Visualization of cross-client data distribution in different domains, performing t-SNE dimensionality reduction on retrieved instructions through various augmentation strategies. We randomly sample 10k in-domain samples as background while randomly sampling 5k samples from the cross-client augmented dataset for different instruction augmentation methods for comparison.